OpenStudy (anonymous):

please multiply 10^1/2 by 100^1/6.

OpenStudy (anonymous):

10^(1/6)

OpenStudy (anonymous):

\[\sqrt{10}\sqrt[6]{100} = 6.8129\]

OpenStudy (anonymous):

or or 10^{5/6}

OpenStudy (anonymous):

hello by implies divide not multiply

OpenStudy (anonymous):

garg90, how did you get 10^{5/6} ?

OpenStudy (anonymous):

my computer don't let me post I try many time

OpenStudy (anonymous):

\[10^{\frac{5}{6}}\]is correct answer

OpenStudy (anonymous):

\[10^{\frac{1}{2}}\times \sqrt[6]{10^{2}}\]

OpenStudy (anonymous):

\[10^{\frac{1}{2}}\times(10^{2})^{\frac{1}{6}}\]

OpenStudy (anonymous):

\[10^{\frac{1}{2}}\times10^{\frac{2}{6}}\]

OpenStudy (anonymous):

\[10^{\frac{1}{2}+\frac{2}{6}}\]

OpenStudy (anonymous):

\[10^{\frac{5}{6}}\]

OpenStudy (anonymous):

ohhh! i get it! thank you so much! :))

OpenStudy (anonymous):

I know now , the computer want post short , can't post long

OpenStudy (anonymous):

thank